OpenStudy (anonymous):

Two dice are rolled. If the total is 2, then player A receives 4 points. If the total is 8, then player B receives 3 points. Find the expected value for each player.

OpenStudy (zarkon):

what are you stuck on?

OpenStudy (anonymous):

i don't know how to set up the problem.

OpenStudy (anonymous):

If the total of the dice is 2, that means that one die is a 1 and the other die is also a 1. The probability of that happening is 1/36 (from 1/6 x 1/6). The expected value is the sum of the probabilities of each occurence times its value. So, for player A, you have (1/36)(4) + (35/36)(0) = 1/9 This is assuming that player A gets no points for any other rolls. Are you with me so far?

OpenStudy (anonymous):

yes! loud and clear :)

OpenStudy (anonymous):

For player B, you have to start by seeing what the probability of rolling an 8 is. You will have the following combinations on the dice: 2,6 3,5 4,4 5,3 6,2 which is 5 rolls out of 36 so the P(8) = 5/36 And player B's expected value is : (5/36)(3) + (31/36)(0) = 5/12 Again, assuming that player B gets no points for any other rolls, similar to the assumption for player A. Is this making sense to you?

OpenStudy (anonymous):

Notice that for player A that 1/36 + 35/36 = 1 and that exhausts all possibilities. Similarly for player B, 5/36 + 31/36 = 1

OpenStudy (anonymous):

thank you so much tcarroll010!!!! i understood everything!!!

OpenStudy (anonymous):

So, the learning points aer that the expected value is the sum of the individual probabilites (times their values) for all possible outcomes.

OpenStudy (anonymous):

And you're quite welcome! Nice working with you!